OpenStudy (anonymous):

HELPPPPPP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! If ∠m ≅ ∠p, m∠m = 9x - 2, and m∠p = 4x + 18, how many degrees are in ∠p?

OpenStudy (anonymous):

you got a picture? i think there are too many m's here

OpenStudy (anonymous):

it might simply mean "solve \(9x-2=4x+18\)"

OpenStudy (anonymous):

No picture. That is the exact question.

OpenStudy (anonymous):

maybe it has \(\large m∠M = 9x - 2\) ?

OpenStudy (anonymous):

no matter, i am pretty sure the real question is to solve \[\large9x-2=4x+18\] do you know how to do it?

OpenStudy (anonymous):

No I do not.

OpenStudy (anonymous):

9x -4x = 18 +2 5x = 20 x =4

OpenStudy (anonymous):

subtract \(4x\) from both sides to get \[\large 5x-2=18\]

OpenStudy (anonymous):

then add 2 to both sides to get \[\large 5x=20\] and finally divide by 5 to get \(x=4\)

OpenStudy (anonymous):

Thats not a answer choice.

OpenStudy (anonymous):

that is because we solved for \(x\) not for \(p\)

OpenStudy (anonymous):

\(p=4x+18\) and \(x=4\) so that means \[p=4\times 4+18\]

OpenStudy (anonymous):

i hope \(34\) is a choice

OpenStudy (anonymous):

that is very confusing satellite73